OpenStudy (anonymous):

Given: f(x)=3x+2 g(x)=x/(x-3) Find (FoG)(x) and list the domain.

hero (hero):

f(g(x)) = 3(x/(x-3))+2

hero (hero):

domain x cannot equal 3

OpenStudy (anonymous):

and that's as simplified as it gets right.

hero (hero):

h(x) = f(g(x)) = 3x/(x-3) + 2

OpenStudy (anonymous):

so when i write it, it would look like this? \[\left(\begin{matrix}3x \\ x-3\end{matrix}\right)+2\]

hero (hero):

\[h(x) =f(g(x)) = \frac{3x}{x-3}+2\]

OpenStudy (anonymous):

THANKS!